My homework problem states "You deposit $2000 into an account that pays for 2% annual interest compounded quarterly. How long will it take for the balance to reach $2500."
using the formula, A = P (1 + r/n)^nt.
I get that I'm solving for t and so far I have 2500 = 2000(1.05)^4t but from then on I'm stuck.
My question is how would I solve this? (You dont have to solve the problem for me, an example would be fine.) Thanks in advance!

Answer by stanbon(75887)

My homework problem states "You deposit $2000 into an account that pays for 2% annual interest compounded quarterly. How long will it take for the balance to reach $2500."
using the formula, A = P (1 + r/n)^nt.
2500 = 2000(1 + (0.02/4))^(4t)
Divide both sides by 2000 to tget:
1.25 = 1.005^4t
Take the log of both sides and solve for "t":
(4t)log(1.005) = log(1.25)
4t = log(1.25)/log(1.005)
4t = 44.74
t = 11.0673
Rounding up: t= 12 years
-------------------------------------
Cheers,
Stan H.